Two compounds with capacity factors of 5.00 and 5.10 are separated on a column with a plate height of 155 µm. What length of column is required to give a resolution of 1.00? How about resolution of 2.00?
Answer #1

if R is resolution N is number of plate K is capacity factor

R=\frac{1}{4}N^{-1/2}(\frac{K_{B}-K_{A}}{1+K_{B}})

1=\frac{1}{4}N^{-1/2}(\frac{5.1-5}{1+5.1})

N = 59536

hight of coloumn = plate hight * number of plate

= .155 m *59536

=9228.080 m

if R= 2 then Number of plate (N) = 238144

hight of coloumn = plate hight * number of plate

=0.155m * 238144

=36912.32 m